# Special Lagrangians and Bridgeland stable objects beyond geometric stability conditions: the product case

### v1

@@ Abstract
where $\mathfrak{s}_{n}$ denotes the product stop. Our main result establishes the existence of a family of stability conditions on this category that has the “stable implies special Lagrangian” property.
@@
    If an object $\mathbb{L}$ is $\sigma_{\mathbf{a},c}$-stable, then it admits a special Lagrangian representative (up to isotopy) in $(\mathbb{C}^{\times})^{n}$ with respect to the holomorphic $n$-form
@@
Recent work by Stoppa [STO25a, STO25b] also investigates the link between Bridgeland stability and the special Lagrangian condition in related contexts. Stoppa’s approach utilizes the theory of deformed Hermitian–Yang–Mills equations [CY21] and primarily addresses stability conditions in the large complex structure limit.
@@
Acknowledgment. The author would like to thank Fabian Haiden, Hansol Hong, Yat-Hin Marco Suen, and Peng Zhou for helpful conversations and correspondences. The author was partially supported by the National Natural Science Foundation of China (Grant No. 12401053).
@@
    If an object $\mathbb{L}$ is $\sigma_{\mathbf{a},c}$-stable, then it admits a special Lagrangian representative (up to isotopy) in $(\mathbb{C}^{\times})^{n}$ with respect to the holomorphic $n$-form
@@
An object $\mathbb{L}$ is $\sigma_{\mathbf{a},c}$-stable if and only if $\mathrm{Mir}_{n}^{-1}(\mathbb{L})$ is $(\sigma_{a_{1}}\boxtimes\cdots\boxtimes\sigma_{a_{n}})$-stable, which by Theorem 3.1(a), happens if and only if
@@
where each $E_{i}$ is $\sigma_{a_{i}}$-stable. By Corollary 2.6, $\mathrm{Mir}(E_{i})$ admits a speical Lagrangian representative with respect to the holomorphic $1$-form
@@
This product (up to isotopy) is a special Lagrangian submanifold with respect to the holomorphic $n$-form
@@
since each factor has constant phase. This concludes the proof. ∎

### v2

@@ Abstract
where $(\mathbb{C}^{\times})^{n}$ is equipped with the split symplectic form
$\omega_{n}\coloneqq\pi_{1}^{*}\omega_{\mathbb{C}^{\times}}+\cdots+\pi_{n}^{*}\omega_{\mathbb{C}^{\times}},$
with $\omega_{\mathbb{C}^{\times}}$ denoting the canonical exact symplectic form on the cotangent bundle $T^{*}S^{1}\cong\mathbb{C}^{\times}$, and $\mathfrak{s}_{n}$ denoting the product stop [GPS24]. Our main result establishes the existence of a family of stability conditions on this category satisfying the property that every stable object is special Lagrangian.
@@
    If an object $\mathbb{L}$ is $\sigma_{\mathbf{a},c}$-stable, then it admits a special Lagrangian representative (up to isotopy) in $((\mathbb{C}^{\times})^{n},\omega_{n})$ with respect to the holomorphic $n$-form
@@
Recent work by Stoppa [STO25a, STO25b] also investigates the link between Bridgeland stability and the special Lagrangian condition in related contexts. Stoppa’s approach utilizes the theory of deformed Hermitian–Yang–Mills equations [CY21] and primarily addresses stability conditions near the large complex structure limit.
@@
Remark 1.8.
We conclude the Introduction with some further questions worth investigating:
  1. (i)
    The converse of Theorem 1.3, namely, whether every special Lagrangian submanifold of $((\mathbb{C}^{\times})^{n},\omega_{n})$ with respect to the holomorphic $n$-form
    is $\sigma_{\mathbf{a},c}$-stable, is not known. A similar difficulty (for different reasons) also appears in a recent work of Stoppa [STO25b], where under suitable conditions, Bridgeland stability implies the existence of a deformed Hermitian–Yang–Mills connection (mirror to special Lagrangian condition near the large volume limit), but the converse is not known.
    In our setting, the simplest testing examples would be to take any $a\in\mathbb{C}$ giving an algebraic stability condition on $H^{0}\mathrm{Tw}(\mathcal{W}(\mathbb{C}^{\times},\mathfrak{s}))\cong\mathrm{D^{b}}(\mathbb{P}^{1})$, and investigate special Lagrangian submanifolds of $((\mathbb{C}^{\times})^{2},\omega_{2})$ with respect to the holomorphic $2$-form
    $\exp\left(z_{1}+z_{2}+\frac{e^{a}}{z_{1}}+\frac{e^{a}}{z_{2}}\right)\frac{\mathrm{d}z_{1}}{z_{1}}\frac{\mathrm{d}z_{2}}{z_{2}}.$
  2. (ii)
    In Remark 4.4, we note that for product-type stability conditions on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ arising from the product of two geometric stability conditions on $\mathbb{P}^{1}$, there are stable objects which are not isomorphic to $E_{1}\boxtimes E_{2}$ for some $E_{1},E_{2}\in\mathrm{D^{b}}(\mathbb{P}^{1})$, unlike those constructed in Theorem 1.3. It would be interesting to see whether the mirror of such an object (which is not a product of two Lagrangians) is special Lagrangian with respect to, say,
    $\exp\left(z_{1}+z_{2}+\frac{e^{b}}{z_{1}}+\frac{e^{b}}{z_{2}}\right)\frac{\mathrm{d}z_{1}}{z_{1}}\frac{\mathrm{d}z_{2}}{z_{2}},$
    where $b\in\mathbb{C}$ gives a geometric stability condition on $H^{0}\mathrm{Tw}(\mathcal{W}(\mathbb{C}^{\times},\mathfrak{s}))\cong\mathrm{D^{b}}(\mathbb{P}^{1})$.
  3. (iii)
    Heuristics in Remark 1.5 suggest an embedding
    $\mathbb{C}^{n}\cong(\mathrm{Stab}\mathrm{D^{b}}(\mathbb{P}^{1})/\mathbb{C})^{n}\hookrightarrow\mathrm{Stab}\mathrm{D^{b}}((\mathbb{P}^{1})^{n})/\mathbb{C}$
    whose image consists of product-type stability conditions as in Definition 1.6. The case $n=2$ is proved in Theorem 1.7. For $n\geq 3$, one would need to show the stability of line bundles $\mathcal{O}(k_{1},\ldots,k_{n})$ with respect to the geometric stability conditions.
  4. (iv)
    More generally, it is natural to ask whether there always exists an embedding
    $(\mathrm{Stab}\mathrm{D^{b}}(X_{1})/\mathbb{C})\times(\mathrm{Stab}\mathrm{D^{b}}(X_{2})/\mathbb{C})\hookrightarrow\mathrm{Stab}\mathrm{D^{b}}(X_{1}\times X_{2})/\mathbb{C}$
    given by product-type stability conditions satisfying the conditions in Definition 1.6, although the symplectic geometric heuristic in Remark 1.5 would no longer be valid.
Acknowledgment. The author would like to thank Fabian Haiden, Hansol Hong, Yu-Shen Lin, Yat-Hin Marco Suen, and Peng Zhou for helpful conversations and correspondences. The author would especially like to thank Jacopo Stoppa for explaining his work and for valuable discussions. The author was partially supported by the National Natural Science Foundation of China (Grant No. 12401053).
@@
    If an object $\mathbb{L}$ is $\sigma_{\mathbf{a},c}$-stable, then it admits a special Lagrangian representative (up to isotopy) in $((\mathbb{C}^{\times})^{n},\omega_{n})$ with respect to the holomorphic $n$-form
@@
An object $\mathbb{L}$ is $\sigma_{\mathbf{a},c}$-stable if and only if $\mathrm{Mir}_{n}^{-1}(\mathbb{L})$ is $(\sigma_{a_{1}}\boxtimes\cdots\boxtimes\sigma_{a_{n}})$-stable, which by Theorem 3.1(a) holds if and only if
@@
where each $E_{i}$ is $\sigma_{a_{i}}$-stable. By Corollary 2.6, $\mathrm{Mir}(E_{i})$ admits a special Lagrangian representative with respect to the holomorphic $1$-form
@@
This is a Lagrangian submanifold of $((\mathbb{C}^{\times})^{n},\omega_{n})$ because the symplectic form $\omega_{n}$ is split, and it is a special Lagrangian submanifold with respect to the holomorphic $n$-form
@@
$=\exp(c)\cdot\prod_{i=1}^{n}\left(\exp\left(z_{i}+\frac{e^{a_{i}}}{z_{i}}\right)\frac{\mathrm{d}z_{i}}{z_{i}}\right),$
since each factor $\mathrm{Mir}(E_{i})$ is of constant phase with respect to $\exp\left(z+\frac{e^{a_{i}}}{z}\right)\frac{\mathrm{d}z}{z}$. This completes the proof. ∎